# On finite group scheme-theoretical categories, II

### v1

@@ 2.8. Principal homogeneous spaces
$\mathscr{O}(G)\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G)\to\mathscr{O}(G)\otimes_{k}\mathscr{O}(H),\,\,\,{\rm f}\otimes\tilde{{\rm f}}\mapsto({\rm f}\otimes 1)\mu_{G\times H}^{\sharp}(\tilde{{\rm f}}),$
@@
[…]
@@
    ${\rm Coh}^{(H,1)}(G)\xrightarrow{\cong}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,(S,\rho)\mapsto\left(S^{(H,1)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta_{\psi}\right)$
    (see 2.21), whose inverse is given by
    ${\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\cong}{\rm Coh}^{(H,1)}(G),\,\,(S,\rho)\mapsto\left(S^{(H,\psi)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta\right).$
@@
    $\pi^{*}:{\rm Coh}(G/H)\xrightarrow{\cong}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,\,U\mapsto\left(U\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G),\rho^{\psi}_{U}\right),$
    whose inverse is given by
    $\pi_{*}^{(H,\psi)}:{\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\cong}{\rm Coh}(G/H),\,\,\,(S,\rho)\mapsto\pi^{(H,\psi)}_{*}S.$
@@
The cleaving map also induces an $\mathscr{O}(H\backslash G)$-linear map
@@
Let $\Phi\in C^{3}(G,\mathbb{G}_{m})$ be any $3$-cochain, and $\xi\in C^{2}(H,\mathbb{G}_{m})$ such that $d\xi=\iota^{\sharp\otimes 3}(\Phi)$. Assume further that $\Theta\in C^{2}(G,\mathbb{G}_{m})$ satisfies $\iota^{\sharp\otimes 2}(\Theta)=\xi$ and $\left(\iota^{\sharp\otimes 2}\otimes{\rm id}\right)(\Phi\cdot d\Theta^{-1})=1$. Then
@@ 2.9. Twisting by the inverse $2$-cocycle
(2.31) ${\rm Corep}_{k}(H,\psi)\xrightarrow{\cong}{\rm Corep}(H,\psi^{-1})_{k},\,\,\,(V,\ell)\mapsto(V,\widetilde{\ell}),$
where ~$\widetilde{\ell}:V\to V\otimes\mathscr{O}(H)_{\psi^{-1}}$, $v\mapsto v^{0}\otimes{\rm S}(v^{-1}Q_{\psi}^{-1})$, is an equivalence of abelian categories.
@@ 2.10. Module categories over ${\rm Coh}(G,\omega)$
    $\mathscr{M}(1,1)\cong{\rm Coh}(G,\omega)$ is the regular ${\rm Coh}(G,\omega)$-module.
@@
    $\mathscr{M}(G,1)\cong{\rm Vec}$ is the usual fiber functor on ${\rm Coh}(G)$.
@@
    The abelian equivalence $\mathscr{M}(H,\psi)\cong{\rm Comod}(\mathscr{O}(H)_{\psi})_{{\rm Coh}(G,\omega)}$ given in Proposition 2.12 is a ${\rm Coh}(G,\omega)$-equivalence.
@@
$S_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}S_{\bar{1}},\,\,\,{\rm and}\,\,\,P_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}P_{\bar{1}}$
@@
$\mathscr{N}(H,\psi):={}^{(H,\psi)}{\rm Coh}(G,\omega)\cong{\rm Comod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})$
@@ 2.11. Module categories over ${\mathscr{C}}(G,\omega,H,\psi)$
${\mathscr{C}}\left(G,\omega,1,1\right)\cong{\rm Coh}(G,\omega),\,\,\,{\rm and}\,\,\,{\mathscr{C}}\left(G,1,G,1\right)\cong{\rm Rep}_{k}(G).$
@@
    $\mathscr{M}\left((H,\psi),(K,\eta)\right)\cong{\rm Bicomod}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}(H)_{\psi},\mathscr{O}(K)_{\eta}\right).$
@@
    ${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\cong\mathscr{M}\left((H,\psi),(K,\eta)\right).$
@@
${\mathscr{C}}(G,\omega,H,\psi)\cong{\rm Bicomod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})^{{\rm rev}},\,\,\,{\rm and}$
${\mathscr{C}}(G,\omega,H,\psi)\cong{\mathscr{M}}\left((H,\psi),(H,\psi)\right)^{{\rm rev}}.$
@@ 3. Associated biequivariant sheaves
(3.1) $j_{1}:B\xrightarrow{b\mapsto(b,1)}B\times C,\,\,\,{\rm and}\,\,\,j_{2}:C\xrightarrow{c\mapsto(1,c)}B\times C$
@@
(3.2) $\partial_{1}:A\xrightarrow{1:1}B,\,\,\,{\rm and}\,\,\,\partial_{2}:A\xrightarrow{1:1}C$
@@
Let $A\backslash(B\times C)$ be the quotient scheme with respect to $\mu_{A\times(B\times C)}$ (2.25), and
@@
such that, setting $\Theta:=\Psi^{-1}{\rm W}$, we obtain
@@
If $\Phi=1$, we can take ${\rm W}=1$. In this case, we have $\Theta=\Psi^{-1}\in Z^{2}(B\times C,\mathbb{G}_{m})$, so $\partial^{\sharp\otimes 2}\left(\Theta\right)\in Z^{2}(A,\mathbb{G}_{m})$ is automatic. ∎
@@
Consider also the actions of $\mathscr{O}(A\backslash(B\times C))$ on $\mathscr{O}(B)$ and $\mathscr{O}(C)$ via $j_{1}^{\sharp}$ and $j_{2}^{\sharp}$, respectively. Finally, let
@@
Let ${\rm Corep}_{k}(A,\xi^{-1})$ denote the category of finite dimensional left $\mathscr{O}(A)_{\xi^{-1}}$-comodules. Our goal in this section is to construct explicit equivalences of abelian categories
[…]
3.1. The equivalence $\mathscr{B}\cong{\rm Corep}_{k}(A,\xi^{-1})$
@@
(3.17) $\lambda_{\beta}:=\beta\cdot(j_{1}^{\sharp}\otimes{\rm id})\Delta_{\mathscr{O}(B\times C)},\,\,\,{\rm and}\,\,\,\rho_{\gamma}:=\gamma\cdot({\rm id}\otimes j_{2}^{\sharp})\Delta_{\mathscr{O}(B\times C)},$
@@
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(B\times C,\Phi)\cong{\rm Vec}$ with the unique simple object being $\left(\mathscr{O}(B\times C),\lambda_{\beta},\rho_{\gamma}\right)$.
@@
    $(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2})\subset(S,\lambda_{2})$ in ${\rm Corep}_{k}(A,\xi^{-1})$.
@@
${\rm t}^{*}:{\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{\cong}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right)$
(see (3.17), and (3.18) below), whose inverse functor is given by
${\rm t}_{*}^{((B,\beta),(C,\gamma),\lambda_{1},\rho)}:\mathscr{B}\xrightarrow{\cong}{\rm Corep}_{k}(A,\xi^{-1}),\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
@@
Let $(V,\ell)\in{\rm Corep}_{k}(A,\xi^{-1})$, and write $\ell(v)=\sum v^{-1}\otimes v^{0}$. Consider the free $\mathscr{O}(B\times C)$-module ${\rm t}^{*}V=\mathscr{O}(B\times C)\otimes_{k}V$, and map
@@
${\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right).$
Conversely, take any $(S,\lambda,\rho)\in\mathscr{B}$, and consider the sheaf ${\rm t}_{*}S$ (that is, the underlying vector space of $S$). Then by Lemma 3.2(4), we have $\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right)\in{\rm Corep}_{k}(A,\xi^{-1})$. Thus, we have a functor
$\mathscr{B}\xrightarrow{}{\rm Corep}_{k}(A,\xi^{-1}),\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
@@
If $A$, $B$ and $C$ are any affine group schemes, then Theorem 3.3 and its proof hold after replacing ${\rm Coh}$ by ${\rm Coh}_{{\rm f}}$ (see [G1]). ∎
3.2. The equivalence $\mathscr{A}\cong\mathscr{B}$
Recall that we are assuming that
@@ 3.2. The equivalence $\mathscr{A}\cong\mathscr{B}$
The functor ${\rm p}^{*}:{\rm Coh}(A\backslash(B\times C))\to{\rm Coh}(B\times C)$ (see §2.1 and (3.4) lifts to an equivalence of abelian categories
${\rm p}^{*}:\mathscr{A}\xrightarrow{\cong}\mathscr{B},\,\,\,(S,\lambda,\rho)\mapsto\left(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S,\lambda^{*},\rho^{*}\right)$
@@
${\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}:\mathscr{B}\xrightarrow{\cong}\mathscr{A},\,\,\,(S,\lambda,\rho)\mapsto({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho).$
@@
If $A$, $B$ and $C$ are affine group schemes such that the quotient $A\backslash(B\times C)$ is affine, then Theorem 3.6 and its proof hold after replacing ${\rm Coh}$ by ${\rm Coh}_{{\rm f}}$ (see [G1]).
3.3. The first equivalence ${\rm Corep}_{k}(A,\xi^{-1})\cong\mathscr{A}$

### v2

@@ 2.9. Principal homogeneous spaces
[…]
@@
(2.19) $\sigma:\mathscr{O}(H)^{\otimes\,2}\to\mathscr{O}(G/H),\,\,\,f\otimes\tilde{f}\mapsto\mathfrak{c}(f_{1})\mathfrak{c}(\tilde{f}_{1})\mathfrak{c}^{-1}(f_{1}\tilde{f}_{2}),$
@@
    ${\rm Coh}^{(H,1)}(G)\xrightarrow{\simeq}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,(S,\rho)\mapsto\left(S^{(H,1)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta_{\psi}\right),$
    whose inverse is given by
    ${\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\simeq}{\rm Coh}^{(H,1)}(G),\,\,(S,\rho)\mapsto\left(S^{(H,\psi)}\otimes_{k}\mathscr{O}(H),{\rm id}\otimes\Delta\right)$
    (see (2.24)).
@@
    $\pi^{*}:{\rm Coh}(G/H)\xrightarrow{\simeq}{\rm Coh}^{(H,\psi)}(G,\omega),\,\,\,U\mapsto\left(U\otimes_{\mathscr{O}(G/H)}\mathscr{O}(G),\rho^{\psi}_{U}\right)$
    (see (2.23)), whose inverse is given by
    $\pi_{*}^{(H,\psi)}:{\rm Coh}^{(H,\psi)}(G,\omega)\xrightarrow{\simeq}{\rm Coh}(G/H),\,\,\,(S,\rho)\mapsto\pi^{(H,\psi)}_{*}S.$
@@
Remark 2.17.
Note that $\omega$ plays no role in Theorem 2.16. ∎
@@
[…]
@@
Let $\Phi\in C^{3}(G,\mathbb{G}_{m})$ be any $3$-cochain, and $\xi\in C^{2}(H,\mathbb{G}_{m})$, such that $d\xi=\iota^{\sharp\otimes 3}(\Phi)$. Assume further that $\Theta\in C^{2}(G,\mathbb{G}_{m})$ satisfies $\iota^{\sharp\otimes 2}(\Theta)=\xi$ and $\left(\iota^{\sharp\otimes 2}\otimes{\rm id}\right)(\Phi\cdot d\Theta^{-1})=1$. Then
@@ 2.10. Twisting by the inverse $2$-cocycle
[…]
where $\widetilde{\ell}:V\to V\otimes\mathscr{O}(H)_{\psi^{-1}}$, $v\mapsto v^{0}\otimes{\rm S}(v^{-1}Q_{\psi}^{-1})$, is an equivalence of abelian categories.
@@ 2.11. Module categories over ${\rm Coh}(G,\omega)$
    $\mathscr{M}(1,1)\simeq{\rm Coh}(G,\omega)$ is the regular ${\rm Coh}(G,\omega)$-module.
@@
    $\mathscr{M}(G,1)\simeq{\rm Vec}$ is the usual fiber functor on ${\rm Coh}(G)$.
@@
    The abelian equivalence $\mathscr{M}(H,\psi)\simeq{\rm Comod}(\mathscr{O}(H)_{\psi})_{{\rm Coh}(G,\omega)}$ given in Proposition 2.14 is a ${\rm Coh}(G,\omega)$-equivalence.
@@
$S_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}S_{\bar{1}}\,\,\,{\rm and}\,\,\,P_{\bar{g}}\cong\delta_{g}\otimes^{{\mathscr{M}}}P_{\bar{1}}$
@@
$\mathscr{N}(H,\psi):={}^{(H,\psi)}{\rm Coh}(G,\omega)\simeq{\rm Comod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})$
@@ 2.12. Module categories over ${\mathscr{C}}(G,\omega,H,\psi)$
${\mathscr{C}}\left(G,\omega,1,1\right)\simeq{\rm Coh}(G,\omega)\,\,\,{\rm and}\,\,\,{\mathscr{C}}\left(G,1,G,1\right)\simeq{\rm Rep}_{k}(G).$
@@
    $\mathscr{M}\left((H,\psi),(K,\eta)\right)\simeq{\rm Bicomod}_{{\rm Coh}(G,\omega)}\left(\mathscr{O}(H)_{\psi},\mathscr{O}(K)_{\eta}\right).$
@@
    ${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\simeq\mathscr{M}\left((H,\psi),(K,\eta)\right).$
@@
${\mathscr{C}}(G,\omega,H,\psi)\simeq{\rm Bicomod}_{{\rm Coh}(G,\omega)}(\mathscr{O}(H)_{\psi})^{{\rm rev}},$
${\mathscr{C}}(G,\omega,H,\psi)\simeq{\mathscr{M}}\left((H,\psi),(H,\psi)\right)^{{\rm rev}}.$
@@ 3. Associated biequivariant sheaves
(3.1) $j_{1}:B\xrightarrow{b\mapsto(b,1)}B\times C\,\,\,{\rm and}\,\,\,j_{2}:C\xrightarrow{c\mapsto(1,c)}B\times C$
@@
(3.2) $\partial_{1}:A\xrightarrow{1:1}B\,\,\,{\rm and}\,\,\,\partial_{2}:A\xrightarrow{1:1}C$
@@
Let $A\backslash(B\times C)$ be the quotient scheme with respect to $\mu_{A\times(B\times C)}$ (2.27), and let
@@
such that, setting $\Theta:=\Psi^{-1}{\rm W}$, we have
@@
If $\Phi=1$, then we can take ${\rm W}=1$. In this case, we have $\Theta=\Psi^{-1}\in Z^{2}(B\times C,\mathbb{G}_{m})$, so $\xi=\partial^{\sharp\otimes 2}\left(\Theta\right)\in Z^{2}(A,\mathbb{G}_{m})$ is automatic. ∎
@@
Consider also the actions of $\mathscr{O}(A\backslash(B\times C))$ on $\mathscr{O}(B)$ and $\mathscr{O}(C)$ via $({\rm p}j_{1})^{\sharp}$ and $({\rm p}j_{2})^{\sharp}$, respectively (see (3.1), (3.4)). Finally, let
@@
Let ${\rm Rep}(A,\xi^{-1})_{k}={\rm Corep}_{k}(\mathscr{O}(A)_{\xi^{-1}})$ be the category of finite dimensional left $\mathscr{O}(A)_{\xi^{-1}}$-comodules. Our goal in this section is to construct explicit equivalences of abelian categories
${\rm Rep}(A,\xi^{-1})_{k}$$\mathscr{B}$$\hskip 1.42271pt\mathscr{A}.$$\scriptstyle{{\rm t}^{*}}$$\scriptstyle{{\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\eta))}}$$\scriptstyle{{\rm t}_{*}^{((B,\beta),(C,\gamma))}}$$\scriptstyle{{}^{(A,\xi^{-1})}{\rm p}_{*}}$$\scriptstyle{{\rm p}^{*}}$$\scriptstyle{{\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}}$
3.1. The equivalence $\mathscr{B}\simeq{\rm Rep}(A,\xi^{-1})_{k}$
@@
(3.17) $\lambda_{\beta}:=(\beta\otimes 1)\cdot(j_{1}^{\sharp}\otimes{\rm id})\Delta_{\mathscr{O}(B\times C)},\,\,\,{\rm and}\,\,\,\rho_{\gamma}:=(1\otimes\gamma)\cdot({\rm id}\otimes j_{2}^{\sharp})\Delta_{\mathscr{O}(B\times C)},$
@@
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(B\times C,\Phi)\simeq{\rm Vec}$ with the unique simple object being $\left(\mathscr{O}(B\times C),\lambda_{\beta},\rho_{\gamma}\right)$.
@@
    $(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2})\subset(S,\lambda_{2})$ in ${\rm Rep}(A,\xi^{-1})_{k}$.
@@
${\rm t}^{*}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right)$
(see (3.17)-(3.18) below), whose inverse functor is given by
${\rm t}_{*}^{((B,\beta),(C,\gamma),\lambda_{1},\rho)}:\mathscr{B}\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
@@
Let $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, and write $\ell(v)=\sum v^{-1}\otimes v^{0}$. Consider the free $\mathscr{O}(B\times C)$-module ${\rm t}^{*}V=\mathscr{O}(B\times C)\otimes_{k}V$, and map
@@
${\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{}\mathscr{B},\,\,\,(V,\ell)\mapsto\left(\mathscr{O}(B\times C)\otimes_{k}V,\lambda^{(\ell,\beta)},\rho_{\gamma}\otimes_{k}{\rm id}_{V}\right).$
Conversely, take any $(S,\lambda,\rho)\in\mathscr{B}$, and consider the sheaf ${\rm t}_{*}S$ (that is, the underlying vector space of $S$). Then by Lemma 3.2(4), we have $\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right)\in{\rm Rep}(A,\xi^{-1})_{k}$. Thus, we have a functor
$\mathscr{B}\xrightarrow{}{\rm Rep}(A,\xi^{-1})_{k},\,\,\,(S,\lambda,\rho)\mapsto\left(S^{((B,\beta),(C,\gamma),\lambda_{1},\rho)},\lambda_{2}\right).$
@@
If $A$, $B$ and $C$ are any affine group schemes, then Theorem 3.3 and its proof hold after replacing $\rm Coh$} by ${\rm Coh}_{{\rm f}}$ (see [G1, Definition 3.2]). ∎
3.2. The equivalence $\mathscr{A}\simeq\mathscr{B}$
Recall that we assume that
@@ 3.2. The equivalence $\mathscr{A}\simeq\mathscr{B}$
The functor ${\rm p}^{*}:{\rm Coh}(A\backslash(B\times C))\to{\rm Coh}(B\times C)$ (see §2.2 and (3.4)) lifts to an equivalence of abelian categories
${\rm p}^{*}:\mathscr{A}\xrightarrow{\simeq}\mathscr{B},\,\,\,(S,\lambda,\rho)\mapsto\left(\mathscr{O}(B\times C)\otimes_{\mathscr{O}(A\backslash(B\times C))}S,\lambda^{*},\rho^{*}\right)$
@@
${\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}:\mathscr{B}\xrightarrow{\simeq}\mathscr{A},\,\,\,(S,\lambda,\rho)\mapsto({\rm p}^{((A,\xi^{-1}),\lambda_{2})}_{*}S,\lambda_{1},\rho).$
@@
If $A$, $B$ and $C$ are affine group schemes such that the quotient $A\backslash(B\times C)$ is affine, then Theorem 3.6 and its proof hold after replacing $\rm Coh$} by ${\rm Coh}_{{\rm f}}$ (see [G1, Definition 3.2]).
3.3. The first equivalence ${\rm Rep}(A,\xi^{-1})_{k}\simeq\mathscr{A}$